you roll a number cube 30 times. About how many times would you expect to roll a 4
Answer by stanbon(75887)

you roll a number cube 30 times. About how many times would you expect to roll a 4
---
P(roll 4) = 1/6
---
Expected # of 4 in 30 rolls = (1/6)*30 = 5
===================
Cheers,
Stan H.
===================